OpenStudy (muscrat123):

Maria looked at the architectural plan of a room with four walls in which the walls meet each other at right angles. The length of one wall in the plan is 13 inches. The length of the diagonal of the floor of the room in the plan is approximately 15.26 inches. Is the room in the shape of a square? Explain how you determined your answer. Show all your work @Sepeario

OpenStudy (sepeario):

This question is essentially the same as the other. If you actually understood the concept, you would be able to do this. Can you try to attempt this question?

OpenStudy (muscrat123):

i did

OpenStudy (muscrat123):

lemme draw it

OpenStudy (muscrat123):

\[15.26^{2} = 13^{2} +b^{2} ~~~~~232.87 = 169 + b ^{2}~~~~~232.87=169 + 63.87~~~~~15.26=13^{2}+7.99^{2}\]

OpenStudy (muscrat123):

*13^2+7.99^2

OpenStudy (muscrat123):

@Sepeario

OpenStudy (muscrat123):

i dont think it is a square because a square has all equal sides. this means that the missing length should be 13, not 7.99

OpenStudy (muscrat123):

OOOH YAY

OpenStudy (sepeario):

Very good. You have proved it correctly.

OpenStudy (muscrat123):

sorry. i didnt realize this until after i thought about whay i just did all those calculations. lol

OpenStudy (muscrat123):

*why

OpenStudy (muscrat123):

but wait

OpenStudy (muscrat123):

"four walls in which the walls meet each other at right angles" -and- it doesn't ask for a missing length is it still not a square?

OpenStudy (muscrat123):

@Sepeario

OpenStudy (sepeario):

It is still isn't a square.

OpenStudy (muscrat123):

y not?

OpenStudy (muscrat123):

?

OpenStudy (muscrat123):

@Sepeario

OpenStudy (muscrat123):

@Hero @jim_thompson5910

OpenStudy (muscrat123):

@uybuyvf

OpenStudy (muscrat123):

@heretohelpalways